OpenStudy (anonymous):

Change 7.2 grams to milligrams.

OpenStudy (godlovesme):

Multiply 7.2 by 1000 :)

OpenStudy (danjs):

\[\frac{ 7.2 g }{ 1 }\frac{ 1 mg }{ 10^{-3} g } = .....mg\]

OpenStudy (danjs):

multiply by ratios of 1, and keep track of how units cancel

OpenStudy (danjs):

always refer your conversions to the base unit first , for mass (g)

OpenStudy (anonymous):

ok

OpenStudy (danjs):

For instance , kg to mg \[\frac{ 7.2 kg }{ 1 }\frac{ 10^3 g }{ 1 kg }\frac{ 1 mg }{ 10^{-3} g} = .....mg\]

OpenStudy (danjs):

notice, kg and g cancel out and you are left with 7.2x10^6 mg

OpenStudy (danjs):

That is what i mean by converting to the base unit , for mass (g) first

OpenStudy (anonymous):

I 100% agree with you @DanJS

OpenStudy (danjs):

So all you have to remember, is how each prefix relates to the base unit, that is it

OpenStudy (danjs):

Same for if you are switching systems. To figure if you need to multiply or divide, set up the fractions so that units cancel in your favor,, IE \[\frac{ 5280 ft }{ 1 mi } \]or \[\frac{ 1 mi }{ 5280 ft }\]depending on how your units need to cancel

OpenStudy (anonymous):

agreed !!!!! you deserve a medal :)

OpenStudy (danjs):



OpenStudy (danjs):

Multiply by ratios of 1. easy as that

OpenStudy (danjs):

wory about the numbers at the end, units are the most important